The number of ways in which letters of the word ‘KEJRIWAL’ be arranged in such a way that the vowels are always in alphabetical order is P. Then P/960 is

Grade:11

1 Answers

BNB
7 years ago
KEJRIWAL in this word there are 3 vowels so we select the all 3 vowels and arrange all remaining letters i.e p=8C3×5!=6720 now, p/960=6720/960=7.